OpenStudy (anonymous):

it takes John 12 seconds to run from point A to point B and Mary 15 seconds to run the same distance. if John runs from A to B and Mary runs from B to A, starting at the same time, how many seconds later will they meet?

OpenStudy (anonymous):

t = d/r d/12 + d/15 = d/t t = 20/3 seconds or 6.67 seconds rounded to the nearest 1/100.